What is the price of a

6​-year,

7.7%

coupon​ rate,

$1,000

face value bond that pays interest annually if the yield to maturity on similar bonds is

6.8%​?

The price of the bond is

​$___

​ (Round to the nearest​ cent.)
Answer #1

Annual coupon=1000*7.7%=77

Hence current price=Annual coupon*Present value of annuity factor(6.8%,6)+1000*Present value of discounting factor(6.8%,6)

=77*4.79611244+1000*0.673864354

=$1043.17(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=77[1-(1.068)^-6]/0.068

=77*4.79611244

2.Present value of discounting factor=1000/1.068^6

=1000*0.673864354